# No Dimensional Sampling Coresets for Classification

### v1

@@ 3.3 Proof of Theorem 10
$M_{1}^{n}$ $=\operatorname*{\mathbb{E}}_{x_{1:m}\sim q}\operatorname*{\mathbb{E}}_{\sigma_{1:m}\sim\{-1,1\}}\sup_{2^{n-1}\leq\|w\|_{H}\leq 2^{n}}\alpha(w)\left|\frac{1}{m}\sum_{i=1}^{m}\sigma_{i}\frac{\bar{\phi}(\boldsymbol{K}(w,x_{i}))}{s(x_{i})}\right|$
(Since $\alpha(w)\leq\frac{kS}{2^{2(n-1)}}$) $\leq\frac{kS}{m4^{n-1}}\operatorname*{\mathbb{E}}_{x_{1:m}\sim q}\operatorname*{\mathbb{E}}_{\sigma_{1:m}\sim\{-1,1\}}\sup_{2^{n-1}\leq\|w\|_{H}\leq 2^{n}}\left|\sum_{i=1}^{m}\sigma_{i}\frac{\bar{\phi}(\boldsymbol{K}(w,x_{i}))}{s(x_{i})}\right|$
(Cor. 16 with $\varepsilon_{i}=\frac{1}{s(x_{i})}$) $\leq\frac{2LkS}{m4^{n-1}}\operatorname*{\mathbb{E}}_{x_{1:m}\sim q}\operatorname*{\mathbb{E}}_{\sigma_{1:m}\sim\{-1,1\}}\sup_{2^{n-1}\leq\|w\|_{H}\leq 2^{n}}\left|\sum_{i=1}^{m}\sigma_{i}\frac{\boldsymbol{K}(w,x_{i})}{s(x_{i})}\right|$
$=\frac{2LkS}{m4^{n-1}}\operatorname*{\mathbb{E}}_{x_{1:m}\sim q}\operatorname*{\mathbb{E}}_{\sigma_{1:m}\sim\{-1,1\}}\sup_{2^{n-1}\leq\|w\|_{H}\leq 2^{n}}\left|\left\langle\boldsymbol{K}(w,\cdot),\sum_{i=1}^{m}\sigma_{i}\frac{\boldsymbol{K}(x_{i},\cdot)}{s(x_{i})}\right\rangle_{H}\right|$
$\leq\frac{2LkS}{m4^{n-1}}\operatorname*{\mathbb{E}}_{x_{1:m}\sim q}\operatorname*{\mathbb{E}}_{\sigma_{1:m}\sim\{-1,1\}}\sup_{2^{n-1}\leq\|w\|_{H}\leq 2^{n}}\|w\|_{H}\left\|\sum_{i=1}^{m}\sigma_{i}\frac{\boldsymbol{K}(x_{i},\cdot)}{s(x_{i})}\right\|_{H}$
$\leq\frac{LkS}{m2^{n-3}}\operatorname*{\mathbb{E}}_{x_{1:m}\sim q}\operatorname*{\mathbb{E}}_{\sigma_{1:m}\sim\{-1,1\}}\left\|\sum_{i=1}^{m}\sigma_{i}\frac{\boldsymbol{K}(x_{i},\cdot)}{s(x_{i})}\right\|_{H}$
@@
which completes the proof.
@@
Theorem 18
[Theorem 10, Restated] For a well behaved probability measure $P$ (see Definition 8), if $s(\cdot):\mathcal{X}\longrightarrow(1,\infty)$ is an upper sensitivity function for $\mathcal{L}^{H}_{\phi,k}$ with total sensitivity $S$ and $m\geq\frac{2S}{\varepsilon^{2}}\left(8C^{2}+S\log\frac{2}{\delta})\right)$, then, with probability at least $1-\delta$, any $s$-sensitivity sample $x_{1},\ldots,x_{m}$ from $\mathcal{X}$ with weights $u_{i}=\frac{S}{ms(x_{i})}$ provides an $\varepsilon$-coreset for $(\mathcal{X},\ P,\ \mathcal{X},\ \ell^{H}_{\phi,k})$, where
@@
Proof  The proof immediately follows from Theorem 14 and Lemma 17.
@@
4 Coresets for Monotonic Functions
Let $H$ be a reproducing kernel Hilbert space of real-valued functions on $\mathcal{X}$ with kernel $\boldsymbol{K}\colon\mathcal{X}\times\mathcal{X}\longrightarrow\mathbb{R}$ and $P$ a probability measure over $\mathcal{X}$. For a non-increasing $L$-Lipschitz function $\phi\colon\mathbb{R}\longrightarrow(0,\infty)$, we remind that $\ell^{H}_{\phi,k}:\mathcal{X}\times\mathcal{X}\longrightarrow\mathbb{R}$ where $\ell^{H}_{\phi,k}(x,w)=\phi(\boldsymbol{K}(x,w))+\frac{1}{k}\|w\|_{H}^{2}.$ An $\varepsilon$-coreset for $(\mathcal{X},P,\mathcal{L}^{H}_{\phi,k})$, called $\varepsilon$-coreset for a monotonic function $\phi$ with respect to $H$, is a set $X=\{x_{1},\ldots,x_{m}\}\subseteq\mathcal{X}$ accompanied by a measure (or weight function) $\nu$ such that
@@ 4 Coresets for Monotonic Functions
Proof  For $t\geq 0$, we should check the inequality
@@
[…]
@@
Proof  Due to similarity, we only prove the lemma for $\ell^{H}_{\phi,k}$. First, note that as $\phi$ is non-increasing, $\beta(\alpha)\geq 1$ for each $\alpha$. Consider arbitrary $x,y,w\in\mathcal{X}$ with $\|x\|_{H},\|y\|_{H}\leq B_{1},\|w\|_{H}\leq B_{2}$. In the following, we deal with the two different cases $\boldsymbol{K}(x,w)\leq 0$ and $\boldsymbol{K}(x,w)>0$ separately. If $\boldsymbol{K}(x,w)\leq 0$, then
$\phi(\boldsymbol{K}(y,w))$ $=\phi(\langle y,w\rangle_{H})$
@@
$\leq\frac{M}{\phi(0)}\phi(\boldsymbol{K}(x,w))$ (since $\boldsymbol{K}(x,w)\leq 0$ and $\phi$ is non-increasing)
@@
Now, assume that $\boldsymbol{K}(x,w)>0$. Similarly,
$\phi(\boldsymbol{K}(y,w))\leq M\leq\frac{M}{\phi(0)}\phi(-\boldsymbol{K}(x,w))\leq\frac{M}{\phi(0)}\phi(-\|x\|_{H}\|w\|_{H}).$
@@
$\ell^{H}_{\phi,k}(y,w)=\phi(\boldsymbol{K}(y,w))+\frac{\|w\|_{H}^{2}}{k}$ $\leq\frac{M}{\phi(0)}\left(\phi(-\|x\|_{H}\|w\|_{H})+\frac{\|w\|_{H}^{2}}{k}\right)$
(Property 13) $\leq\frac{M}{\phi(0)}\beta(\|x\|_{H})\left(\phi(\|x\|_{H}\|w\|_{H})+\frac{\|w\|_{H}^{2}}{k}\right)$
(Cauchy-Schwarts Inequality and $\phi$ is non-increasing) $\leq\frac{M}{\phi(0)}\beta(\|x\|_{H})\underbrace{\left(\phi(\boldsymbol{K}(x,w))+\frac{\|w\|_{H}^{2}}{k}\right)}_{=\ell^{H}_{\phi,k}(x,w)},$
@@
Combining Equations (14) and (15), we obtain the desired inequality.
@@
Proof  For a $y\in\mathcal{X}$, we have
@@
completing the proof.
It is worth noting that the provided upper sensitivity function $s(y)=\frac{1}{\int_{\mathcal{X}}\gamma(\|x\|){\rm d}P}$ remains constant, irrespective of $y\in\mathcal{X}$. In a specific scenario where $\gamma$ is lower bounded by $L$, the function $s(y)=\frac{1}{L}$ serves as an upper sensitivity function, possessing a total value of $S=\frac{1}{L}$. It is important to note that when $\phi$ and $k$ grantee Property (16) for $\ell=\ell^{H}_{\phi,k}$ (like what is established in Lemma 20), then Lemma 21 provides a methodology to calculate the upper sensitivity function and its total value for $\mathcal{L}^{H}_{\phi,k}$. This enables us to obtain $s$-sensitivity samples from $\mathcal{X}$, making Theorem 18 (and Theorem 3) applicable in this context.
4.1 Coreset for Sigmoid function
@@
Using this result, they established the first item of Theorem 6 in which, as a drawback, $k$ should be sufficiently large contingent on the given data points $\mathcal{X}$. This dependency comes from the role of $k_{\alpha}$ in Equation (17). In the subsequent lemma, we eliminate this dependency.
Lemma 22
@@
Proof  One can verify that $\sigma:\mathbb{R}\longrightarrow(0,1)$ is a positive decreasing function and $\sigma^{\prime}(t)=-\sigma(t)\sigma(-t)$. For a fixed $\alpha>0$, if we set $t=\alpha z$ and $K=\alpha^{2}k$,
@@
Note that for $\alpha=0$ or $z=0$, the inequality is valid as well.
Lemma 23
Assume that $H$ is a reproducing kernel Hilbert space of real-valued functions on $\mathcal{X}$ with kernel $\boldsymbol{K}\colon\mathcal{X}\times\mathcal{X}\longrightarrow\mathbb{R}$ and $P$ is a probability measure over $\mathcal{X}$ such that $\operatorname*{\mathbb{E}}\limits_{x\sim P}\|x\|_{H}^{2}\leq E^{2}_{1}.$ Then $s(x)=60+32kE_{1}^{2}$ is an upper sensitivity function for $(\mathcal{X},\ P,\ \ell^{H}_{\sigma,k})$ and $(\mathcal{X},\ P,\ \bar{\ell}^{H}_{\sigma,k})$.
Proof  Given $\operatorname*{\mathbb{E}}_{x\sim P}(\|x\|_{H}^{2})\leq E^{2}_{1}$, Markov’s inequality implies that $\operatorname*{\mathbb{P}}_{x\sim P}(\|x\|_{H}\geq 2E^{2}_{1})\leq\frac{1}{2}$ and consequently $\operatorname*{\mathbb{P}}_{x\sim P}(\|x\|_{H}\leq 2E^{2}_{1})\geq\frac{1}{2}$ which will be used later to derive inequality marked by in (*)(*)( * ). Since $\sigma$ is universally bounded by $1$. Lemma 22 indicates that Lemma 20 is applicable with
@@
and for all $x,y,w\in\mathcal{X}$ and $M=1$. Using Lemma 21 with $\gamma(\|x\|_{H})=\frac{\sigma(0)}{M\beta_{\sigma}(\|x\|_{H})}=\frac{1}{2\beta_{\sigma}(\|x\|_{H})}$, we conclude that $s(y)=\frac{2}{\int_{\mathbb{R}^{d}}\frac{1}{\beta_{\sigma}(\|x\|_{H})}{\rm d}P}$ is an upper sensitivity function for $(\mathcal{X},\ P,\ \mathcal{X},\ \ell^{H}_{\sigma,k})$ and $(\mathcal{X},\ P,\ \mathcal{X},\ \bar{\ell}^{H}_{\sigma,k})$. Note that
@@
This concludes $s(x)=60+32kE_{1}^{2}$ is an upper sensitivity function for $(\mathcal{X},P,\ell^{H}_{\sigma,k})$ and $(\mathcal{X},P,\bar{\ell}^{H}_{\sigma,k})$.
Now, we are at a point to put Theorem 10 into action.
Theorem 24 (No Dimensional Sigmoid Coreset)
Assume that $H$ is a reproducing kernel Hilbert space of real-valued functions on $\mathcal{X}$ with kernel $\boldsymbol{K}\colon\mathcal{X}\times\mathcal{X}\longrightarrow\mathbb{R}$, $P$ is a probability measure over $\mathcal{X}$ such that $\operatorname*{\mathbb{E}}\limits_{x\sim p}\|x\|_{H}^{2}\leq E^{2}_{1}$, $S=60+32kE_{1}^{2}$, and $C=\left(2E_{1}+\frac{1}{2}\right)\max\left(4E_{1}k,\frac{2}{5}\right)+8kE_{1}+1.$ For $m\geq\frac{2S}{\varepsilon^{2}}(8C^{2}+S\log\frac{2}{\delta})$, any iid sample $x_{1},\ldots,x_{m}$ from $P$ with weights $u_{i}=\frac{1}{m}$ provides an $\varepsilon$-coreset for $(\mathcal{X},P,\mathcal{X},\ell^{H}_{\sigma,k})$ (respectively for $(\mathcal{X},P,\mathcal{X},\bar{\ell}^{H}_{\sigma,k})$ ) with probability at least $1-\delta$.
Proof  Lemma 23 indicates that $s(x)=60+32kE^{2}_{1}$ is an upper sensitivity function. Given that it is a constant function, $s$-sensitivity sampling is equivalent to sampling according to $P$. Notice that $\sigma$ is a $1$-Lipschitz function. As $\log(\cdot)$ is a concave function, using Jensen’s inequality, we have $\operatorname*{\mathbb{E}}\left(\log\sigma\left(\frac{\|x_{i}\|_{H}}{2E_{1}}\right)\right)\leq\log\operatorname*{\mathbb{E}}\left(\sigma\left(\frac{\|x_{i}\|_{H}}{2E_{1}}\right)\right).$ On the other hand, since $-\log\sigma(t)=\log(1+e^{t})\leq 1+t,$ we have
@@
Hence, $P$ is well behaved (see Definition 8), thereby allowing Theorem 10 to establish the statement.
While Theorem 24 holds for any reproducing kernel Hilbert space meeting the specified criteria, the most intriguing instance arises in Hilbert space $\mathbb{R}^{d}$ equipped with the standard inner product (also known as the dot product) as its kernel. In this scenario, we can also employ Theorem 3, which leads to the subsequent theorem.
Theorem 25
@@
Proof  Because of similarity, we only prove the statement for $(\mathbb{R}^{d},P,\mathbb{R}^{d},\ell_{\sigma,k})$. Given that Lemma 23 asserts the constant function $s(x)=60+32kE^{2}_{1}$ as an upper sensitivity function for $\mathcal{L}_{\sigma,k}$, we deduce $\operatorname{\bf Ranges}(\mathcal{T}_{\mathcal{L}_{\sigma,k}},\succ)=\operatorname{\bf Ranges}(\mathcal{L}_{\sigma,k},\succ)$. To clarify, every function in $\mathcal{T}_{\mathcal{L}_{\sigma,k}}$ is a function in $\mathcal{L}_{\sigma,k}$ scaled by a positive constant. This implies ${\rm VCdim}(\operatorname{\bf Ranges}(\mathcal{T}_{\mathcal{L}_{\sigma,k}},\succ))={\rm VCdim}(\operatorname{\bf Ranges}(\mathcal{L}_{\sigma,k},\succ)).$ We would like to remind that
@@
which concludes that $\operatorname{\bf Ranges}(\mathcal{L}_{\sigma,k},\succ)$ only includes half-spaces, empty set, and the whole space $\mathbb{R}^{d}$. Therefore, by Radon’s theorem, ${\rm VCdim}(\operatorname{\bf Ranges}(\mathcal{L}_{\sigma,k},\succ))\leq d+1$ (for a proof, see Lemma 10.3.1 in  (Matousek, 2002)). Leveraging Theorem 3, we derive the statement for $m=O(\frac{S}{\varepsilon^{2}}\left(d\log S+\log\frac{1}{\delta}\right))$, thereby completing the proof.
Theorems 24 and 25 can be considered as an advancement of the first item of Theorem 6. Theorem 6 is applicable when $\mathcal{X}$ is a finite set inside the unit ball, and $m$ is function that is linear in terms of $\log(|\mathcal{X}|)$ and quadratic in terms $d$, whereas in Theorem 24, $\mathcal{X}$ can be infinite, and moreover, the presented $m$ remains constant concerning those parameters (assuming $E_{1}$ is independent of $d$). Moreover, in Theorem 25, $m$ linearly depends on $d$, whereas in Theorem 6, it has a quadratic dependence on $d$.
4.2 Coreset for Logistic Function

### v2

@@ C.2 Contraction Lemma
$M_{1}^{n}$ $=\operatorname*{\mathbb{E}}_{x_{1:m}\sim q}\operatorname*{\mathbb{E}}_{\sigma_{1:m}\sim\{-1,1\}}\sup_{2^{n-1}\leq\|w\|_{H}\leq 2^{n}}\alpha(w)\left|\frac{1}{m}\sum_{i=1}^{m}\sigma_{i}\frac{\bar{\phi}(\mathsf{K}(w,x_{i}))}{s(x_{i})}\right|$
(Since $\alpha(w)\leq\frac{kS}{2^{2(n-1)}}$) $\leq\frac{kS}{m4^{n-1}}\operatorname*{\mathbb{E}}_{x_{1:m}\sim q}\operatorname*{\mathbb{E}}_{\sigma_{1:m}\sim\{-1,1\}}\sup_{2^{n-1}\leq\|w\|_{H}\leq 2^{n}}\left|\sum_{i=1}^{m}\sigma_{i}\frac{\bar{\phi}(\mathsf{K}(w,x_{i}))}{s(x_{i})}\right|$
(Cor. C.4 with $\varepsilon_{i}=\frac{1}{s(x_{i})}$) $\leq\frac{2LkS}{m4^{n-1}}\operatorname*{\mathbb{E}}_{x_{1:m}\sim q}\operatorname*{\mathbb{E}}_{\sigma_{1:m}\sim\{-1,1\}}\sup_{2^{n-1}\leq\|w\|_{H}\leq 2^{n}}\left|\sum_{i=1}^{m}\sigma_{i}\frac{\mathsf{K}(w,x_{i})}{s(x_{i})}\right|$
$=\frac{2LkS}{m4^{n-1}}\operatorname*{\mathbb{E}}_{x_{1:m}\sim q}\operatorname*{\mathbb{E}}_{\sigma_{1:m}\sim\{-1,1\}}\sup_{2^{n-1}\leq\|w\|_{H}\leq 2^{n}}\left|\left\langle\mathsf{K}(w,\cdot),\sum_{i=1}^{m}\sigma_{i}\frac{\mathsf{K}(x_{i},\cdot)}{s(x_{i})}\right\rangle_{H}\right|$
$\leq\frac{2LkS}{m4^{n-1}}\operatorname*{\mathbb{E}}_{x_{1:m}\sim q}\operatorname*{\mathbb{E}}_{\sigma_{1:m}\sim\{-1,1\}}\sup_{2^{n-1}\leq\|w\|_{H}\leq 2^{n}}\|w\|_{H}\left\|\sum_{i=1}^{m}\sigma_{i}\frac{\mathsf{K}(x_{i},\cdot)}{s(x_{i})}\right\|_{H}$
$\leq\frac{LkS}{m2^{n-3}}\operatorname*{\mathbb{E}}_{x_{1:m}\sim q}\operatorname*{\mathbb{E}}_{\sigma_{1:m}\sim\{-1,1\}}\left\|\sum_{i=1}^{m}\sigma_{i}\frac{\mathsf{K}(x_{i},\cdot)}{s(x_{i})}\right\|_{H}$
@@
which completes the proof. ∎
@@
Theorem C.6.
[Theorem A.3, Restated] For a well behaved probability measure $P$ (see Definition A.1), if $s(\cdot):\mathcal{X}\longrightarrow(1,\infty)$ is an upper sensitivity function for $\mathcal{L}^{H}_{\phi,k}$ with total sensitivity $S$ and $m\geq\frac{2S}{\varepsilon^{2}}\left(8C^{2}+S\log\frac{2}{\delta})\right)$, then, with probability at least $1-\delta$, any $s$-sensitivity sample $x_{1},\ldots,x_{m}$ from $\mathcal{X}$ with weights $u_{i}=\frac{S}{ms(x_{i})}$ provides an $\varepsilon$-coreset for $(\mathcal{X},\ P,\ \mathcal{X},\ \ell^{H}_{\phi,k})$, where
@@
Proof.
The proof immediately follows from Theorem C.2 and Lemma C.5. ∎
@@
Appendix D Coresets for Monotonic Functions
Let $H$ be a reproducing kernel Hilbert space of real-valued functions on $\mathcal{X}$ with kernel $\mathsf{K}\colon\mathcal{X}\times\mathcal{X}\longrightarrow\mathbb{R}$ and $P$ a probability measure over $\mathcal{X}$. For a non-increasing $L$-Lipschitz function $\phi\colon\mathbb{R}\longrightarrow(0,\infty)$, we remind that $\ell^{H}_{\phi,k}:\mathcal{X}\times\mathcal{X}\longrightarrow\mathbb{R}$ where $\ell^{H}_{\phi,k}(x,w)=\phi(\mathsf{K}(x,w))+\frac{1}{k}\|w\|_{H}^{2}.$ An $\varepsilon$-coreset for $(\mathcal{X},P,\mathcal{L}^{H}_{\phi,k})$, called $\varepsilon$-coreset for a monotonic function $\phi$ with respect to $H$, is a set $X=\{x_{1},\ldots,x_{m}\}\subseteq\mathcal{X}$ accompanied by a measure (or weight function) $\nu$ such that
@@ Appendix D Coresets for Monotonic Functions
Proof.
For $t\geq 0$, we should check the inequality
@@
$\frac{\phi(0)}{M\beta(\|x\|_{H})}\leq\frac{\ell^{H}_{\phi,k}(x,w)}{\ell^{H}_{\phi,k}(y,w)}\qquad\text{and}\qquad\frac{\phi(0)}{M\beta(\|x\|_{H})}\leq\frac{\bar{\ell}^{H}_{\phi,k}(x,w)}{\bar{\ell}^{H}_{\phi,k}(y,w)}.$
@@
Proof.
Due to similarity, we only prove the lemma for $\ell^{H}_{\phi,k}$. First, note that as $\phi$ is non-increasing, $\beta(\alpha)\geq 1$ for each $\alpha$. Consider arbitrary $x,y,w\in\mathcal{X}$ with $\|x\|_{H},\|y\|_{H}\leq B_{1},\|w\|_{H}\leq B_{2}$. In the following, we deal with the two different cases $\mathsf{K}(x,w)\leq 0$ and $\mathsf{K}(x,w)>0$ separately. If $\mathsf{K}(x,w)\leq 0$, then
$\phi(\mathsf{K}(y,w))$ $=\phi(\langle y,w\rangle_{H})$
@@
$\leq\frac{M}{\phi(0)}\phi(\mathsf{K}(x,w))$ (since $\mathsf{K}(x,w)\leq 0$ and $\phi$ is non-increasing)
@@
Now, assume that $\mathsf{K}(x,w)>0$. Similarly,
$\phi(\mathsf{K}(y,w))\leq M\leq\frac{M}{\phi(0)}\phi(-\mathsf{K}(x,w))\leq\frac{M}{\phi(0)}\phi(-\|x\|_{H}\|w\|_{H}).$
@@
$\ell^{H}_{\phi,k}(y,w)=\phi(\mathsf{K}(y,w))+\frac{\|w\|_{H}^{2}}{k}$ $\leq\frac{M}{\phi(0)}\left(\phi(-\|x\|_{H}\|w\|_{H})+\frac{\|w\|_{H}^{2}}{k}\right)$
(Property 21) $\leq\frac{M}{\phi(0)}\beta(\|x\|_{H})\left(\phi(\|x\|_{H}\|w\|_{H})+\frac{\|w\|_{H}^{2}}{k}\right)$
(Cauchy-Schwarts Inequality and $\phi$ is non-increasing) $\leq\frac{M}{\phi(0)}\beta(\|x\|_{H})\underbrace{\left(\phi(\mathsf{K}(x,w))+\frac{\|w\|_{H}^{2}}{k}\right)}_{=\ell^{H}_{\phi,k}(x,w)},$
@@
Combining Equations (22) and (23), we obtain the desired inequality. ∎
@@
Proof.
For a $y\in\mathcal{X}$, we have
@@
completing the proof. ∎
It is worth noting that the provided upper sensitivity function $s(y)=\frac{1}{\int_{\mathcal{X}}\gamma(\|x\|){\rm d}P}$ remains constant, irrespective of $y\in\mathcal{X}$. In a specific scenario where $\gamma$ is lower bounded by $L$, the function $s(y)=\frac{1}{L}$ serves as an upper sensitivity function, possessing a total value of $S=\frac{1}{L}$. It is important to note that when $\phi$ and $k$ grantee Property (24) for $\ell=\ell^{H}_{\phi,k}$ (like what is established in Lemma D.2), then Lemma D.3 provides a methodology to calculate the upper sensitivity function and its total value for $\mathcal{L}^{H}_{\phi,k}$. This enables us to obtain $s$-sensitivity samples from $\mathcal{X}$, making Theorem C.6 (and Theorem 1.2) applicable in this context.
D.1 Coreset for Sigmoid function
@@
Using this result, they established the first item of Theorem 1.5 in which, as a drawback, $k$ should be sufficiently large contingent on the given data points $\mathcal{X}$. This dependency comes from the role of $k_{\alpha}$ in Equation (25). In the subsequent lemma, we eliminate this dependency.
Lemma D.4.
@@
Proof.
One can verify that $\sigma:\mathbb{R}\longrightarrow(0,1)$ is a positive decreasing function and $\sigma^{\prime}(t)=-\sigma(t)\sigma(-t)$. For a fixed $\alpha>0$, if we set $t=\alpha z$ and $K=\alpha^{2}k$,
@@
Note that for $\alpha=0$ or $z=0$, the inequality is valid as well. ∎
Lemma D.5.
Assume that $H$ is a reproducing kernel Hilbert space of real-valued functions on $\mathcal{X}$ with kernel $\mathsf{K}\colon\mathcal{X}\times\mathcal{X}\longrightarrow\mathbb{R}$ and $P$ is a probability measure over $\mathcal{X}$ such that $\operatorname*{\mathbb{E}}\limits_{x\sim P}\|x\|_{H}^{2}\leq E^{2}_{1}.$ Then $s(x)=60+32kE_{1}^{2}$ is an upper sensitivity function for $(\mathcal{X},\ P,\ \ell^{H}_{\sigma,k})$ and $(\mathcal{X},\ P,\ \bar{\ell}^{H}_{\sigma,k})$.
Proof.
Given $\operatorname*{\mathbb{E}}_{x\sim P}(\|x\|_{H}^{2})\leq E^{2}_{1}$, Markov’s inequality implies that $\operatorname*{\mathbb{P}}_{x\sim P}(\|x\|_{H}\geq 2E^{2}_{1})\leq\frac{1}{2}$ and consequently $\operatorname*{\mathbb{P}}_{x\sim P}(\|x\|_{H}\leq 2E^{2}_{1})\geq\frac{1}{2}$ which will be used later to derive inequality marked by in ()(*)( ∗ ). Since $\sigma$ is universally bounded by $1$. Lemma D.4 indicates that Lemma D.2 is applicable with
@@
and for all $x,y,w\in\mathcal{X}$ and $M=1$. Using Lemma D.3 with $\gamma(\|x\|_{H})=\frac{\sigma(0)}{M\beta_{\sigma}(\|x\|_{H})}=\frac{1}{2\beta_{\sigma}(\|x\|_{H})}$, we conclude that $s(y)=\frac{2}{\int_{\mathbb{R}^{d}}\frac{1}{\beta_{\sigma}(\|x\|_{H})}{\rm d}P}$ is an upper sensitivity function for $(\mathcal{X},\ P,\ \mathcal{X},\ \ell^{H}_{\sigma,k})$ and $(\mathcal{X},\ P,\ \mathcal{X},\ \bar{\ell}^{H}_{\sigma,k})$. Note that
@@
This concludes $s(x)=60+32kE_{1}^{2}$ is an upper sensitivity function for $(\mathcal{X},P,\ell^{H}_{\sigma,k})$ and $(\mathcal{X},P,\bar{\ell}^{H}_{\sigma,k})$. ∎
Now, we are at a point to put Theorem A.3 into action.
Theorem D.6 (No Dimensional Sigmoid Coreset).
Assume that $H$ is a reproducing kernel Hilbert space of real-valued functions on $\mathcal{X}$ with kernel $\mathsf{K}\colon\mathcal{X}\times\mathcal{X}\longrightarrow\mathbb{R}$, $P$ is a probability measure over $\mathcal{X}$ such that $\operatorname*{\mathbb{E}}\limits_{x\sim p}\|x\|_{H}^{2}\leq E^{2}_{1}$, $S=60+32kE_{1}^{2}$, and $C=\left(2E_{1}+\frac{1}{2}\right)\max\left(4E_{1}k,\frac{2}{5}\right)+8kE_{1}+1.$ For $m\geq\frac{2S}{\varepsilon^{2}}(8C^{2}+S\log\frac{2}{\delta})$, any iid sample $x_{1},\ldots,x_{m}$ from $P$ with weights $u_{i}=\frac{1}{m}$ provides an $\varepsilon$-coreset for $(\mathcal{X},P,\mathcal{X},\ell^{H}_{\sigma,k})$ (respectively for $(\mathcal{X},P,\mathcal{X},\bar{\ell}^{H}_{\sigma,k})$ ) with probability at least $1-\delta$.
Proof.
Lemma D.5 indicates that $s(x)=60+32kE^{2}_{1}$ is an upper sensitivity function. Given that it is a constant function, $s$-sensitivity sampling is equivalent to sampling according to $P$. Notice that $\sigma$ is a $1$-Lipschitz function. As $\log(\cdot)$ is a concave function, using Jensen’s inequality, we have $\operatorname*{\mathbb{E}}\left(\log\sigma\left(\frac{\|x_{i}\|_{H}}{2E_{1}}\right)\right)\leq\log\operatorname*{\mathbb{E}}\left(\sigma\left(\frac{\|x_{i}\|_{H}}{2E_{1}}\right)\right).$ On the other hand, since $-\log\sigma(t)=\log(1+e^{t})\leq 1+t,$ we have
@@
Hence, $P$ is well behaved (see Definition A.1), thereby allowing Theorem A.3 to establish the statement. ∎
While Theorem D.6 holds for any reproducing kernel Hilbert space meeting the specified criteria, the most intriguing instance arises in Hilbert space $\mathbb{R}^{d}$ equipped with the standard inner product (also known as the dot product) as its kernel. In this scenario, we can also employ Theorem 1.2, which leads to the subsequent theorem.
Theorem D.7.
@@
Proof.
Because of similarity, we only prove the statement for $(\mathbb{R}^{d},P,\mathbb{R}^{d},\ell_{\sigma,k})$. Given that Lemma D.5 asserts the constant function $s(x)=60+32kE^{2}_{1}$ as an upper sensitivity function for $\mathcal{L}_{\sigma,k}$, we deduce $\operatorname{\bf Ranges}(\mathcal{T}_{\mathcal{L}_{\sigma,k}},\succ)=\operatorname{\bf Ranges}(\mathcal{L}_{\sigma,k},\succ)$. To clarify, every function in $\mathcal{T}_{\mathcal{L}_{\sigma,k}}$ is a function in $\mathcal{L}_{\sigma,k}$ scaled by a positive constant. This implies ${\rm VCdim}(\operatorname{\bf Ranges}(\mathcal{T}_{\mathcal{L}_{\sigma,k}},\succ))={\rm VCdim}(\operatorname{\bf Ranges}(\mathcal{L}_{\sigma,k},\succ)).$ We would like to remind that
@@
which concludes that $\operatorname{\bf Ranges}(\mathcal{L}_{\sigma,k},\succ)$ only includes half-spaces, empty set, and the whole space $\mathbb{R}^{d}$. Therefore, by Radon’s theorem, ${\rm VCdim}(\operatorname{\bf Ranges}(\mathcal{L}_{\sigma,k},\succ))\leq d+1$ (for a proof, see Lemma 10.3.1 in  (Matousek, 2002)). Leveraging Theorem 1.2, we derive the statement for $m=O(\frac{S}{\varepsilon^{2}}\left(d\log S+\log\frac{1}{\delta}\right))$, thereby completing the proof. ∎
Theorems D.6 and D.7 can be considered as an advancement of the first item of Theorem 1.5. Theorem 1.5 is applicable when $\mathcal{X}$ is a finite set inside the unit ball, and $m$ is function that is linear in terms of $\log(|\mathcal{X}|)$ and quadratic in terms $d$, whereas in Theorem D.6, $\mathcal{X}$ can be infinite, and moreover, the presented $m$ remains constant concerning those parameters (assuming $E_{1}$ is independent of $d$). Moreover, in Theorem D.7, $m$ linearly depends on $d$, whereas in Theorem 1.5, it has a quadratic dependence on $d$.
D.2 Coreset for Logistic Function